2. A study was conducted to test the effect of a new drug (Clofibrate) on reducing the coronary death compared to placebo. Patients were randomized into two groups at 1:2 ratio. Of 1065 subjects in the treatment group, 194 died at 5 years, whereas 523 of 2695 subjects in the placebo group died at 5 years. Answer these questions: a) what is the study design? b) generate a 2 by 2 table, c) calculate the crude mortality rates for both groups, d) calculate relative risk of dying for treatment group as compared to placebo, e) interpret your result in one sentence. (20 points - 4 each)
Answer #1

a) The experimental study design is used in this research study. In this, the subjects randomized into two groups. One group received a new drug(treatment group) and other groups receive a placebo(control group) and the outcomes are compared and studied.

b) Generate 2 X 2 table.

Outcome

Death

(5 years)

Total
Exposed to a new drug(Clofibrate) 871 194 1065
Placebo 2442 523 2965
Total 3313 717 4030

C) Crude death rate = No. of death / Average population X 100000

Crude death rate for treatment group = 194/ 1065 X 100,000

= 18216 death/ 100,000 population.

The crude Death rate for placebo group = 523/2965 X 100,000

= 17639 death/ 100,000

d) Calculate Relative Risk RR = A/(A+B) / C/(C+D).

A = No. of death who underwent the treatment of new drug = 194.

B = Subtract A from the total = 1065 - 194 = 871.

C = No. of death in the placebo group = 523.

D= Subtract C from the total = 2965 -523 = 2442.

Relative risk = (194/1065) / (523/2965)

= (0.182) / (0.176)

= 1.0340

The relative risk of occuring coronary death with a new drug is 1.0340.

e) The relative risk is greater than one. It shows that there is a greater risk of coronary death in the treatment of new drug compared to placebo.